1.  A hollow metal sphere has an outer diameter of 4cm and an inside diameter of 3.85 cm

a.      What is the thickness in cm?

b.     What is the outer volume cm3?

c.      What is the inner volume cm3?

d.     What is the volume of the metal sphere cm3?


1
Expert's answer
2021-02-25T17:44:52-0500

a)


"s=0.5(4-3.85)=0.075\ cm"

b)


"V_{out}=\frac{4}{3}\pi(4)^3=268\ cm^3"

c)


"V_{in}=\frac{4}{3}\pi(3.85)^3=239\ cm^3"

d)


"V=268-239=29\ cm^3"